# “What If My Face Gets Scanned Without Consent”: Understanding Older Adults’ Experiences with Biometric Payment

### v1

@@ 6. Conclusion
Our work makes the first attempt to explore older adults’ perceptions and practices regarding biometric payment. To achieve this, we conducted semi-structured interviews with 22 older adults in China. Our findings reveal that convenience, security, social influence, and facilitating conditions are key motivators for using biometric payment. However, older adults also expressed various concerns, including lack of control, security and privacy risks (e.g., unauthorized biometric data collection and biometric data leakage), adaptability challenges, and lack of usage knowledge. To balance these benefits and risks, many older adults actively adopt protective measures, such as limiting biometric data disclosure. They also expressed a desire for future improvements, including the integration of lightweight, personalized, and context-aware cognitive confirmation mechanisms, as well as more comprehensive educational support. Based on these findings, we discuss three key tensions: the balance between ease of use and security and privacy concerns, the password-free benefit versus lack of control, and the gap between perceived understanding and practical reality. We also position biometric payment within the broader contexts of digital payment and biometric technology and provide actionable implications to enhance user control, empower autonomy, and facilitate protective measures.

### v2

@@ 6. Conclusion
Our work makes the first attempt to explore older adults’ perceptions and practices regarding biometric payment. To achieve this, we conducted semi-structured interviews with 22 older adults in China. Our findings reveal that convenience, security, social influence, and facilitating conditions are key motivators for using biometric payment. However, older adults expressed various concerns, including lack of control, security and privacy risks (e.g., unauthorized biometric data collection and biometric data leakage), adaptability challenges, and lack of usage knowledge. To balance these benefits and risks, many older adults actively adopt protective measures, such as limiting biometric data disclosure. They also identified possible improvements, including the integration of lightweight, personalized, and context-aware cognitive confirmation mechanisms and more comprehensive educational support. Based on these findings, we discuss three key tensions: the balance between ease of use and security and privacy concerns, the password-free benefit versus lack of control, and the gap between perceived understanding and practical reality. We also position biometric payment within the broader contexts of digital payment and biometric technology and provide recommendations to enhance user control, empower autonomy, and facilitate protective measures.
Acknowledgements.
The research is partially funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2092 CASA - 390781972, NSFC grant 62432008, RGC RIF grant R6021-20, RGC TRS grant T43-513/23N-2, RGC CRF grant C6015-23G, NSFC/RGC grant CRS_HKUST601/24, and RGC GRF grants 16207922, 16207423 and 16203824. We thank Lin Kyi for her valuable feedback on our early drafts.